AnakKrakatau1883:

You get a 50% on a test, which is 30/60 points. If you get an extra 20 points, what percentage will you have on the test? Show your work.

Astro:

Well to start we have 30 points, what is 30+20?

AnakKrakatau1883:

50.

Astro:

Exactly! What do we do next?

AnakKrakatau1883:

We have that left over 50%. I presume we convert it into a decimal.

Astro:

Correct. And that would be?

AnakKrakatau1883:

0.5.

AnakKrakatau1883:

Per cent = Per 100.

Astro:

Sorry, I am one strep ahead of myself. That is not what we would do next. What we should do is 50/60

Astro:

@astro wrote:
Sorry, I am one strep ahead of myself. That is not what we would do next. What we should do is 50/60
Step, my bad

AnakKrakatau1883:

0.83 is the answer. Which means it would be 83% But let's say I got 10 extra out of the possible 20. What would that come down to?

Shadow:

You can get a bit lazy about this. Since 30 is 50% of 60, And we're adding 20 points, All we need to do to find the final percentage is add what percent of 20 is of 60. 20/60 = 1/3, so we just add 33.3%

Astro:

@shadow wrote:
You can get a bit lazy about this. Since 30 is 50% of 60, And we're adding 20 points, All we need to do to find the final percentage is add what percent of 20 is of 60. 20/60 = 1/3, so we just add 33.3%
Thanks, Shadow.

Shadow:

Doing 50/60 will be the more widely applicable approach to less pretty numbers. But here you can cut corners because the numbers are pretty.

Astro:

@anakkrakatau1883 wrote:
0.83 is the answer. Which means it would be 83% But let's say I got 10 extra out of the possible 20. What would that come down to?
Well we can do what we were just doing but with 10 instead of 20

Astro:

So we would have 30+10=40 40/60=.66*100=%

Astro:

@astro wrote:
So we would have 30+10=40 40/60=.66*100=%
^66%

Astro:

You can also make it very simple by just doing 4/6 which should be easy...66%

AnakKrakatau1883:

@astro wrote:
@astro wrote:
So we would have 30+10=40 40/60=.66*100=%
^66%
Thanks, Astro!

Astro:

@anakkrakatau1883 wrote:
@astro wrote:
@astro wrote:
So we would have 30+10=40 40/60=.66*100=%
^66%
Thanks, Astro!
Happy to help and thanks to @Shadow for the back-up

AnakKrakatau1883:

All I know is Biology, World History, and 4 other classes that I'm taking at FLVS.

AnakKrakatau1883:

But I do know meteorology, a bit of chemistry, and some oceanography.